what is the mass of 2.90x10^2 L of a liquid if the density is 1.025 g/cm^3
1
Expert's answer
2013-09-16T09:52:31-0400
2.90x10^2L=2.90*10^5 ml. So, m = 2.90*10^5 ml*1.025 g/cm^3 = 2.9725*10^5 g